OpenStudy (anonymous):

HELP PLEASE!!! In driving the usual route to school a student computes his average speed to be 30km/h. In a hurry to get home that afternoon he wants to average 60km/h for the round trip. What would be his average speed for the return trip over the same route have to be in order to do this????? Support mathematically??

OpenStudy (anonymous):

what grade are you in?

OpenStudy (anonymous):

Y?

OpenStudy (anonymous):

Y??

OpenStudy (anonymous):

so i know how to help you

OpenStudy (anonymous):

11th

OpenStudy (anonymous):

45 i think?!

OpenStudy (anonymous):

The mathematical formula: The average speed for roundtrip=(The average speed to get to school + the average speed to get to your home)/2 keep in mind that the formula for the mean (average) is the sum of all of the numbers in a list divided by the number of items in that list.

OpenStudy (anonymous):

see i was right :D

OpenStudy (anonymous):

lets treat the average speed to get home as our variable x:\[(30+x)\div2=60\]

OpenStudy (anonymous):

can you work the problem out from here on?

OpenStudy (anonymous):

Ok but our book says its impossible but I have to prove that and that's y I'm confused I had your answer originally but the I was told it would have to be infinity for it to work but I need to prove this mathematically

OpenStudy (anonymous):

who are you talking to Lulu212?:D

OpenStudy (anonymous):

Either one of u

OpenStudy (anonymous):

By the way Lorii95 is wrong, if you work out the problem you can see that x = 90: \[(30+x)=60*2\] \[x=120-30\] \[x = 90\]

OpenStudy (anonymous):

Ok thanks